what is the equation of the line that passes through the point (5,11) and is parallel to y= 4/5x+1 ?
Answer by scott8148(6628)

parallel means the slope is the same

use the given point to find the new y-intercept

11 = (4/5)5 + b ___ 11 = 4 + b ___ 7 = b

the new equation is ___ y = (4/5)x + 7